# Reshuffling Resampling Splits Can Improve Generalization of Hyperparameter Optimization

### v1

@@ 1 Introduction
Hyperparameters have been shown to strongly influence the performance of machine learning models (van Rijn & Hutter, 2018; Probst et al., 2019). The primary objective of hyperparameter optimization (HPO; also called tuning) is the identification and selection of a hyperparameter configuration (HPC) that minimizes the estimated generalization error (Feurer & Hutter, 2019; Bischl et al., 2023). Typically, this task is challenged by the absence of a closed-form mathematical description of the objective function, the unavailability of an analytic gradient, and the large cost to evaluate HPCs, categorizing HPO as a noisy, black-box optimization problem. An HPC is evaluated via resampling, such as a holdout split or $M$-fold cross-validation (CV), during tuning.
@@
These resampling splits are usually constructed in a fixed and instantiated manner, i.e., the same training and validation splits are used for the internal evaluation of all configurations. On the one hand, this is an intuitive approach, as it should facilitate a fair comparison between HPCs and reduce the variance in the comparison. [^1] This approach likely originates from the concept of paired statistical tests and the resulting variance reduction, but in our literature search we did not find any references discussing this in the context of HPO. For example, when comparing the performance of two classifiers on one dataset, paired tests are commonly employed that implicitly assume that differences between the performance of classifiers on a given CV fold are comparable (Dietterich, 1998; Nadeau & Bengio, 1999, 2003; Demšar, 2006). On the other hand, such a fixing of train and validation splits might steer the optimization, especially after a substantial budget of evaluations, towards favoring HPCs which are specifically tailored to the chosen splits. Such and related effects, where we "overoptimize" the validation performance without effective reward in improved generalization performance have been sometimes dubbed "overtuning" or "oversearching". For a more detailed discussion of this topic, including related work, see Section 5 and Appendix A. The practice of reshuffling resampling splits during HPO is generally neither discussed in the scientific literature nor HPO software tools. [^2] In Appendix A, we present an overview of how resampling is addressed in tutorials and examples of standard HPO libraries and software. We conclude that usually fixed splits are used or recommended. To the best of our knowledge, only Lévesque (2018) investigated reshuffling train-validation splits for every new HPC. For both holdout and $M$-fold CV using reshuffled resampling splits resulted in, on average, slightly lower generalization error when used in combination with Bayesian optimization (BO, Garnett, 2023) or CMA-ES (Hansen & Ostermeier, 2001) as HPO algorithms. ML-Plan (Mohr et al., 2018) proposes a two-stage procedure. In a first stage, the tool uses planning on hierarchical task networks to find promising machine learning pipelines. In a second step, it reshuffles the data and retrains the most promising candidates from the search step to make a more unbiased selection of the final model. Lastly, in the context of evolutionary optimization, reshuffling was applied after every generation (Larcher & Barbosa, 2022).
@@ 2.1 Problem Statement and Setup
Machine learning (ML) aims to fit a model to data, so that it generalizes well to new observations of the same distribution. Let $\mathcal{D}=\{\bm{Z}_{i}\}_{i=1}^{n}$ be the observed dataset consisting of i.i.d. random variables from a distribution $P$, i.e., in the supervised setting $\bm{Z}_{i}=(\bm{X}_{i},Y_{i})$. [^3] Throughout, we use bold letters to indicate (fixed and random) vectors. Formally, an inducer $g$ configured by an HPC $\bm{\lambda}\in\Lambda$ maps a dataset $\mathcal{D}$ to a model from our hypothesis space $h=g_{\bm{\lambda}}(\mathcal{D})\in\mathcal{H}$. During HPO, we want to find a HPC that minimizes the expected generalization error, i.e., find
@@
and the $M$-fold validation loss is
@@
Typically, the same splits are used for every HPC, so $\mathcal{I}_{m,j}=\mathcal{I}_{m}$ for all $j=1,\dots,J$ and $m=1,\dots,M$. In the following, we investigate how reshuffling train-validation splits (i.e., $\mathcal{I}_{m,j}\neq\mathcal{I}_{m,j^{\prime}}$ for $j\neq j^{\prime}$) affects the HPO problem. The exact definitions of the index sets $\mathcal{I}_{m,j}$ used by popular splitting strategies can be found in Appendix D.
@@ 2.2 How Reshuffling Affects the Loss Surface
Under regularity conditions given in Section B.1, it holds
@@
where $\mathcal{T}$ is a training set of size $n$ and $\bm{Z},\bm{Z}^{\prime}$ are two fresh samples from the same distribution.
The kernel $K$ reflects the (co-)variability of the losses caused by validation samples. The contribution of training samples only has a higher-order effect. The validation scheme enters the distribution through the quantities $\tau_{i,j,M}.$ In what follows, we compute explicit expressions for some popular examples. In all cases, we in fact have
$\tau_{i,j,M}=\begin{cases}\sigma^{2},&i=j\\ \tau^{2}\sigma^{2},&i\neq j.\end{cases}.$ (1)
The parameter $\sigma^{2}$ captures any increase in variance caused by omitting an observation from the validation sets. The parameter $\tau$ is the reshuffling effect. The observed losses $\widehat{\mu}(\bm{\lambda}_{i}),\widehat{\mu}(\bm{\lambda}_{j})$ at distinct HPCs $\bm{\lambda}_{i}\neq\bm{\lambda}_{j}$ become less correlated when $\tau$ is small.
Table 1 shows the parameters $\sigma^{2}$ and $\tau^{2}$ for selected methods, see Appendix D for details.
Table 1: Exemplary parametrizations in Equation (1) for resamplings; see Appendix D for details.
@@
We make the following observations:
@@ 2.3 How Reshuffling Affects HPO Performance
Let $\Lambda\subseteq\{\bm{\lambda}\in\mathds{R}^{d}\colon\|\bm{\lambda}\|\leq 1\}$ with $|\Lambda|=J<\infty$ be the set of hyperparameters. Theorem 2.2 ahead gives a bound on the expected regret $\mathds{E}[\mu(\widehat{\bm{\lambda}})]-\mu(\bm{\lambda}^{*})$. It depends on several quantities characterizing the difficulty of the HPO problem. The constant
@@
can be interpreted as a measure of correlation of the process $\epsilon$. In particular, $\mathsf{Corr}(\epsilon(\bm{\lambda}),\epsilon(\bm{\lambda}^{\prime}))\geq 1-\kappa\|\bm{\lambda}-\bm{\lambda}^{\prime}\|^{2}$. The constant is small when $\epsilon$ is strongly correlated, and large otherwise. Further, define $\eta$ as the minimal number such that any $\eta$-ball contained in $\{\|\bm{\lambda}\|\leq 1\}$ contains at least one element of $\Lambda$. It measures how densely the set of candidate HPCs $\Lambda$ covers set of all possible HPCs. If $\Lambda$ is a deterministic uniform grid, we have about $\eta\approx J^{-1/d}$. Similarly, Lemma C.1 in the Appendix shows that $\eta\lesssim J^{-1/2d}$ when randomly sampling HPCs. Finally, the constant
@@
measures the local curvature of the loss surface $\mu$ at its minimum. Finding an HPC $\bm{\lambda}$ close to the theoretical optimum $\bm{\lambda}^{*}$ is easier when the minimum is more pronounced (large $m$). On the other hand, the regret $\mu(\bm{\lambda})-\mu(\bm{\lambda}^{*})$ is also punishing mistakes more quickly. Defining $\log(x)_{+}=\max\{0,\log(x)\}$, we can now state our main result.
@@
The numeric constants result from several simplifications in a worst-case analysis, which lowers their practical relevance. A qualitative analysis of the bound is still insightful. The bound is increasing in $\sigma$ and $d$, indicating that the HPO problem is harder when there is a lot of noise or there are many parameters to tune. The terms $B(\tau)$ and $A(\tau)$ have conceptual interpretations.
@@
Both $A$ and $B$ are decreasing in the reshuffling parameter $\tau$. There are two regimes. If $\sigma/2m\eta^{2}\leq e$, then $A(\tau)=0$ and reshuffling cannot lead to an improvement of the bound. The term $\sigma/m\eta^{2}$ can be interpreted as noise-to-signal ratio (relative to the grid density). If the signal is much stronger than the noise, the HPO problem is so easy that reshuffling will not help.
Figure 1: Example of decorrelated empirical loss yielding better minimizer.
If on the other hand $\sigma/m\eta^{2}>e$, the terms $A(\tau)$ and $B(\tau)$ enter the bound with opposing signs. This creates tension: reshuffling between HPCs increases $B(\tau)$, which is countered by a decrease in $A(\tau)$. So which scenarios favor reshuffling? When the process $\epsilon$ is strongly correlated, $\kappa$ is small and reshuffling (decreasing $\tau$) incurs a high cost in $B(\tau)$. This is intuitive: When there is strong correlation, the validation loss surface $\widehat{\mu}$ is essentially just a vertical shift of $\mu$. Finding the optimal $\bm{\lambda}$ is then almost as easy as if we would know $\mu$, and decorrelating the surface through reshuffling would make it unnecessarily hard. When $\epsilon$ is less correlated ($\kappa$ large) however, reshuffling does not hurt the term $B(\tau)$ as much, but we can reap all the benefits of increasing $A(\tau)$. Here, the effect of reshuffling can be interpreted as hedging against the catastrophic case where all $\widehat{\mu}(\bm{\lambda})$ close to the optimal $\bm{\lambda}^{*}$ are simultaneously dominated by a region of bad hyperparameters. This is illustrated in Figure 1 on the right.
@@ 3.1 Design
We construct a univariate quadratic loss surface function $\mu:\Lambda\subset\mathbb{R}\mapsto\mathbb{R},\lambda\rightarrow m(\lambda-0.5)^{2}/2$ which we want to minimize. The global minimum is given at $\mu(0.5)=0$. Combined with a kernel for the noise process $\epsilon$ as in Equation (3), this allows us to simulate an objective as observed during HPO by sampling $\widehat{\mu}(\lambda)=\mu(\lambda)+\epsilon(\lambda)$. We use a squared exponential kernel $K(\lambda,\lambda^{\prime})=\sigma_{K}^{2}\exp{(-\kappa(\lambda-\lambda^{\prime})^{2}/2)}$ that is plugged into the covariance kernel of the noise process $\epsilon$ (Equation (3)). The parameters $m$ and $\kappa$ in our simulation setup correspond exactly to the curvature and correlation constants from the previous sections. Recall that Theorem 2.2 states that the effect of reshuffling strongly depends on the curvature $m$ of the loss surface $\mu$ (a larger $m$ implies a stronger curvature) and the constant $\kappa$ as a measure of correlation of the noise $\epsilon$ (a larger $\kappa$ implies weaker correlation). Combined with the possibility to vary $\tau$ in the covariance kernel (3) of $\epsilon$, we can systematically investigate how curvature of the loss surface, correlation of the noise and the extent of reshuffling affect optimization performance. In each simulation run, we simulate the observed objective $\hat{\mu}(\lambda)$, identify the minimizer $\hat{\lambda}=\arg\min_{\lambda\in\Lambda}\hat{\mu}(\lambda)$, and calculate its true risk, $\mu(\hat{\lambda})$. We repeat this process $10000$ times for various combinations of $\tau$, $m$, and $\kappa$.
@@ 4 Benchmark Experiments
In this section, we present benchmark experiments of real-world HPO problems where we investigate the effect of reshuffling resampling splits during HPO. First, we discuss the experimental setup. Second, we present results for HPO using random search (Bergstra & Bengio, 2012). Third, we also show the effect of reshuffling when applied in BO using HEBO (Cowen-Rivers et al., 2022).
Recall that our theoretical insight suggests that 1) Reshuffling might be beneficial during HPO and 2) holdout should be affected the most by reshuffling and other resamplings should only be affected to a lesser extent.
@@ 4.1 Experimental Setup
As benchmark tasks, we use a set of standard HPO problems defined on small- to medium-sized sized tabular datasets for binary classification. We suspect the effect of the resampling variant used and whether resampling is reshuffled to be larger for smaller datasets, where the variance of the validation loss estimator is naturally higher. Further, from a practical perspective, this also ensures computational feasibility given the large number of HPO runs in our experiments. We systematically vary the learning algorithm, optimized performance metric, resampling method, whether the resampling is reshuffled, and the size of the dataset used for training and validation during HPO. Below, we outline the general experimental design and refer to Appendix E for details.
@@
We used a subset of the datasets defined by the AutoML benchmark (Gijsbers et al., 2024), treating these as data generating processes (DGPs, Hothorn et al., 2005). We only considered datasets with less than $100$ features to reduce the required computation time and required the number of observations to be between $10000$ and $1000000$; for further details see Appendix E.1. Our aim was to robustly measure the generalization performance when varying the size $n$, which, as defined in Section 2 denotes the size of the combined data for model selection, so one training and validation set combined. First, we sampled $5000$ data points per dataset for robust assessment of the generalization error; these points are not used during HPO in any way. Then, from the remaining points we sampled tasks with $n\in\{500,1000,5000\}$.
@@
We selected CatBoost (Prokhorenkova et al., 2018) and XGBoost (Chen & Guestrin, 2016) for their state-of-the-art performance on tabular data (Grinsztajn et al., 2022; Borisov et al., 2022; McElfresh et al., 2023; Kohli et al., 2024). Additionally, we included an Elastic Net (Zou & Hastie, 2005) to represent a linear baseline with a smaller search space and a funnel-shaped MLP (Zimmer et al., 2021) as a cost-effective neural network baseline. We provide details regarding training pipelines and search spaces in Appendix E.2.
@@
We conduct a random search with $500$ HPC evaluations for every resampling strategy we described in Table 1, for both fixed and reshuffled splits. We always use 80/20 train-validation splits for holdout and 5-fold CVs, so that training set size (and negative estimation bias) are always the same. Anytime test performance of an HPO run is assessed by re-training the current incumbent (i.e. the best HPC until the current HPO iteration based on validation performance) on all available train and validation data and evaluating its performance on the outer test set. Note we do this for scientific evaluation in this experiment; obviously, this is not possible in practice. Using random search allows us to record various metrics and afterwards simulate optimizing for different ones, specifically, we recorded accuracy, area under the ROC curve (ROC AUC) and logloss.
@@
We also investigated the effect of reshuffling on a state-of-the-art BO variant, namely HEBO (Cowen-Rivers et al., 2022; Eggensperger et al., 2021; Turner et al., 2021). The experimental design was the same as for random search, except for the budget, which we reduced from 500 HPCs to 250 HPCs, and only optimized ROC AUC.

### v2

@@ 1 Introduction
Hyperparameters have been shown to strongly influence the performance of machine learning models (van Rijn & Hutter, 2018; Probst et al., 2019). The primary goal of hyperparameter optimization (HPO; also called tuning) is the identification and selection of a hyperparameter configuration (HPC) that minimizes the estimated generalization error (Feurer & Hutter, 2019; Bischl et al., 2023). Typically, this task is challenged by the absence of a closed-form mathematical description of the objective function, the unavailability of an analytic gradient, and the large cost to evaluate HPCs, categorizing HPO as a noisy, black-box optimization problem. An HPC is evaluated via resampling, such as a holdout split or $M$-fold cross-validation (CV), during tuning.
@@
These resampling splits are usually constructed in a fixed and instantiated manner, i.e., the same training and validation splits are used for the internal evaluation of all configurations. On the one hand, this is an intuitive approach, as it should facilitate a fair comparison between HPCs and reduce the variance in the comparison. [^1] This approach likely originates from the concept of paired statistical tests and the resulting variance reduction, but in our literature search we did not find any references discussing this in the context of HPO. For example, when comparing the performance of two classifiers on one dataset, paired tests are commonly employed that implicitly assume that differences between the performance of classifiers on a given CV fold are comparable (Dietterich, 1998; Nadeau & Bengio, 1999, 2003; Demšar, 2006). On the other hand, such a fixing of train and validation splits might steer the optimization, especially after a substantial budget of evaluations, towards favoring HPCs which are specifically tailored to the chosen splits. Such and related effects, where we "overoptimize" the validation performance without effective reward in improved generalization performance have been sometimes dubbed "overtuning" or "oversearching". For a more detailed discussion of this topic, including related work, see Section 5 and Appendix B. The practice of reshuffling resampling splits during HPO is generally neither discussed in the scientific literature nor HPO software tools. [^2] In Appendix B, we present an overview of how resampling is addressed in tutorials and examples of standard HPO libraries and software. We conclude that usually fixed splits are used or recommended. To the best of our knowledge, only Lévesque (2018) investigated reshuffling train-validation splits for every new HPC. For both holdout and $M$-fold CV using reshuffled resampling splits resulted in, on average, slightly lower generalization error when used in combination with Bayesian optimization (BO, Garnett, 2023) or CMA-ES (Hansen & Ostermeier, 2001) as HPO algorithms. Additionally, reshuffling was used by a solution to the NeurIPS 2006 performance prediction challenge to estimate the final generalization performance (Guyon et al., 2006). Recently, in the context of evolutionary optimization, reshuffling was applied after every generation (Larcher & Barbosa, 2022).
@@ 2.1 Problem Statement and Setup
Machine learning (ML) aims to fit a model to data, so that it generalizes well to new observations of the same distribution. Let $\mathcal{D}=\{\bm{Z}_{i}\}_{i=1}^{n}$ be the observed dataset consisting of i.i.d. random variables from a distribution $P$, i.e., in the supervised setting $\bm{Z}_{i}=(\bm{X}_{i},Y_{i})$. [^3] Throughout, we use bold letters to indicate (fixed and random) vectors., [^4] We provide a notation table for symbols used in the main paper in Table 2 in the appendix. Formally, an inducer $g$ configured by an HPC $\bm{\lambda}\in\Lambda$ maps a dataset $\mathcal{D}$ to a model from our hypothesis space $h=g_{\bm{\lambda}}(\mathcal{D})\in\mathcal{H}$. During HPO, we want to find a HPC that minimizes the expected generalization error, i.e., find
@@
and the $M$-fold validation loss as
@@
Typically, the same splits are used for every HPC, so $\mathcal{I}_{m,j}=\mathcal{I}_{m}$ for all $j=1,\dots,J$ and $m=1,\dots,M$. In the following, we investigate how reshuffling train-validation splits (i.e., $\mathcal{I}_{m,j}\neq\mathcal{I}_{m,j^{\prime}}$ for $j\neq j^{\prime}$) affects the HPO problem.
@@ 2.2 How Reshuffling Affects the Loss Surface
Under regularity conditions stated in Appendix C.1, it holds
@@
where the expectation is taken over a training set $\mathcal{T}$ of size $n$ and two fresh samples $\bm{Z},\bm{Z}^{\prime}$ from the same distribution.
The regularity conditions are rather mild and discussed further in Appendix C.1. The kernel $K$ reflects the (co-)variability of the losses caused by validation samples. The contribution of training samples only has a higher-order effect. The validation scheme enters the distribution through the quantities $\tau_{i,j,M}.$ In what follows, we compute explicit expressions for some popular examples. The following list provides formal definitions for the index sets $\mathcal{I}_{m,j}$.
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
  5. (v)
  6. (vi)
The value of $\tau_{i,j,M}$ for each example is computed explicitly in Appendix E. In all these examples, we in fact have
$\tau_{i,j,M}=\begin{cases}\sigma^{2},&i=j\\ \tau^{2}\sigma^{2},&i\neq j.\end{cases},$ (1)
for some method-dependent parameters $\sigma,\tau$ shown in Table 1. The parameter $\sigma^{2}$ captures any increase in variance caused by omitting an observation from the validation sets. The parameter $\tau$ quantifies a potential decrease in correlation in the loss surface due to reshuffling. More precisely, the observed losses $\widehat{\mu}(\bm{\lambda}_{i}),\widehat{\mu}(\bm{\lambda}_{j})$ at distinct HPCs $\bm{\lambda}_{i}\neq\bm{\lambda}_{j}$ become less correlated when $\tau$ is small. Generally, an increase in variance leads to worse generalization performance. The effect of a correlation decrease is less obvious and is studied in detail in the following section.
Table 1: Exemplary parametrizations in Equation (1) for resamplings; see Appendix E for details.
@@
We make the following observations about the differences between methods in Table 1:
@@ 2.3 How Reshuffling Affects HPO Performance
Let $\Lambda\subseteq\{\bm{\lambda}\in\mathds{R}^{d}\colon\|\bm{\lambda}\|\leq 1\}$ with $|\Lambda|=J<\infty$ be the set of hyperparameters. Theorem 2.2 ahead gives a bound on the expected regret $\mathds{E}[\mu(\widehat{\bm{\lambda}})-\mu(\bm{\lambda}^{*})]$. It depends on several quantities characterizing the difficulty of the HPO problem. The constant
@@
can be interpreted as a measure of correlation of the process $\epsilon$. In particular, $\mathsf{Corr}(\epsilon(\bm{\lambda}),\epsilon(\bm{\lambda}^{\prime}))\geq 1-\kappa\|\bm{\lambda}-\bm{\lambda}^{\prime}\|^{2}$. The constant is small when $\epsilon$ is strongly correlated, and large otherwise. Further, define $\eta$ as the minimal number such that any $\eta$-ball contained in $\{\|\bm{\lambda}\|\leq 1\}$ contains at least one element of $\Lambda$. It measures how densely the set of candidate HPCs $\Lambda$ covers set of all possible HPCs. If $\Lambda$ is a deterministic uniform grid, we have about $\eta\approx J^{-1/d}$. Similarly, Lemma D.1 in the Appendix shows that $\eta\lesssim J^{-1/2d}$ when randomly sampling HPCs. Finally, the constant
@@
measures the local curvature at the minimum of the loss surface $\mu$. Finding an HPC $\bm{\lambda}$ close to the theoretical optimum $\bm{\lambda}^{*}$ is easier when the minimum is more pronounced (large $m$). On the other hand, the regret $\mu(\bm{\lambda})-\mu(\bm{\lambda}^{*})$ is also punishing mistakes more quickly. Defining $\log(x)_{+}=\max\{0,\log(x)\}$, we can now state our main result.
@@
The numeric constants result from several simplifications in a worst-case analysis, which lowers their practical relevance. A qualitative analysis of the bound is still insightful. The bound is increasing in $\sigma$ and $d$, indicating that the HPO problem is harder when there is a lot of noise or there are many parameters to tune. The terms $B(\tau)$ and $A(\tau)$ have conceptual interpretations:
@@
The $B$, as stated, is unbounded, but a closer inspection of the proof shows that it is upper bounded by $\sqrt{\log J}$. This bound is attained only in the unrealistic scenario when the validation losses are essentially uncorrelated across all HPCs. The term $A$ is bounded from below by zero, which is also the worst case because the term enters our regret bound with a negative sign.
Both $A$ and $B$ are decreasing in the reshuffling parameter $\tau$. There are two regimes. If $\sigma/2m\eta^{2}\leq e$, then $A(\tau)=0$ and reshuffling cannot lead to an improvement of the bound. The term $\sigma/m\eta^{2}$ can be interpreted as noise-to-signal ratio (relative to the grid density). If the signal is much stronger than the noise, the HPO problem is so easy that reshuffling will not help. This situation is illustrated in Figure 1(a).
(a) High signal-to-noise ratio
(b) Low signal-to-noise ratio
Figure 1: Example of reshuffled empirical loss yielding a worse (left) and better (right) minimizer.
If on the other hand $\sigma/m\eta^{2}>e$, the terms $A(\tau)$ and $B(\tau)$ enter the bound with opposing signs. This creates tension: reshuffling between HPCs increases $B(\tau)$, which is countered by a decrease in $A(\tau)$. So which scenarios favor reshuffling? When the process $\epsilon$ is strongly correlated, $\kappa$ is small and reshuffling (decreasing $\tau$) incurs a high cost in $B(\tau)$. This is intuitive: When there is strong correlation, the validation loss surface $\widehat{\mu}$ is essentially just a vertical shift of $\mu$. Finding the optimal $\bm{\lambda}$ is then almost as easy as if we would know $\mu$, and decorrelating the surface through reshuffling would make it unnecessarily hard. When $\epsilon$ is less correlated ($\kappa$ large) however, reshuffling does not hurt the term $B(\tau)$ as much, but we can reap all the benefits of increasing $A(\tau)$. Here, the effect of reshuffling can be interpreted as hedging against the catastrophic case where all $\widehat{\mu}(\bm{\lambda})$ close to the optimal $\bm{\lambda}^{*}$ are simultaneously dominated by a region of bad hyperparameters. This is illustrated in Figure 1(b).
@@ 3.1 Design
We construct a univariate quadratic loss surface function $\mu:\Lambda\subset\mathbb{R}\mapsto\mathbb{R},\lambda\rightarrow m(\lambda-0.5)^{2}/2$ which we want to minimize. The global minimum is given at $\mu(0.5)=0$. Combined with a kernel for the noise process $\epsilon$ as in Equation (3), this allows us to simulate an objective as observed during HPO by sampling $\widehat{\mu}(\lambda)=\mu(\lambda)+\epsilon(\lambda)$. We use a squared exponential kernel $K(\lambda,\lambda^{\prime})=\sigma_{K}^{2}\exp{(-\kappa(\lambda-\lambda^{\prime})^{2}/2)}$ that is plugged into the covariance kernel of the noise process $\epsilon$ in Equation (3). The parameters $m$ and $\kappa$ in our simulation setup correspond exactly to the curvature and correlation constants from the previous sections. Recall that Theorem 2.2 states that the effect of reshuffling strongly depends on the curvature $m$ of the loss surface $\mu$ (a larger $m$ implies a stronger curvature) and the constant $\kappa$ as a measure of correlation of the noise $\epsilon$ (a larger $\kappa$ implies weaker correlation). Combined with the possibility to vary $\tau$ in the covariance kernel of $\epsilon$, we can systematically investigate how curvature of the loss surface, correlation of the noise and the extent of reshuffling affect optimization performance. In each simulation run, we simulate the observed objective $\hat{\mu}(\lambda)$, identify the minimizer $\hat{\lambda}=\arg\min_{\lambda\in\Lambda}\hat{\mu}(\lambda)$, and calculate its true risk, $\mu(\hat{\lambda})$. We repeat this process $10000$ times for various combinations of $\tau$, $m$, and $\kappa$.
@@ 4 Benchmark Experiments
In this section, we present benchmark experiments of real-world HPO problems where we investigate the effect of reshuffling resampling splits during HPO. First, we discuss the experimental setup. Second, we present results for HPO using random search (Bergstra & Bengio, 2012). Third, we also show the effect of reshuffling when applied in BO using HEBO (Cowen-Rivers et al., 2022) and SMAC3 (Lindauer et al., 2022). Recall that our theoretical insight suggests that 1) reshuffling might be beneficial during HPO and 2) holdout should be affected the most by reshuffling and other resamplings should only be affected to a lesser extent.
@@ 4.1 Experimental Setup
As benchmark tasks, we use a set of standard HPO problems defined on small- to medium-sized tabular datasets for binary classification. We suspect the effect of the resampling variant used and whether the resampling is reshuffled to be larger for smaller datasets, where the variance of the validation loss estimator is naturally higher. Furthermore, from a practical perspective, this also ensures computational feasibility given the large number of HPO runs in our experiments. We systematically vary the learning algorithm, optimized performance metric, resampling method, whether the resampling is reshuffled, and the size of the dataset used for training and validation during HPO. Below, we outline the general experimental design and refer to Appendix F for details.
@@
We used a subset of the datasets defined by the AutoML benchmark (Gijsbers et al., 2024), treating these as data generating processes (DGPs; Hothorn et al., 2005). We only considered datasets with less than $100$ features to reduce the required computation time and required the number of observations to be between $10000$ and $1000000$; for further details see Appendix F.1. Our aim was to robustly measure the generalization performance when varying the size $n$, which, as defined in Section 2 denotes the size of the combined data for model selection, so one training and validation set combined. First, we sampled $5000$ data points per dataset for robust assessment of the generalization error; these points are not used during HPO in any way. Then, from the remaining points we sampled tasks with $n\in\{500,1000,5000\}$.
@@
We selected CatBoost (Prokhorenkova et al., 2018) and XGBoost (Chen & Guestrin, 2016) for their state-of-the-art performance on tabular data (Grinsztajn et al., 2022; Borisov et al., 2022; McElfresh et al., 2023; Kohli et al., 2024). Additionally, we included an Elastic Net (Zou & Hastie, 2005) to represent a linear baseline with a smaller search space and a funnel-shaped MLP (Zimmer et al., 2021) as a cost-effective neural network baseline. We provide details regarding training pipelines and search spaces in Appendix F.2.
@@
We conduct a random search with $500$ HPC evaluations for every resampling strategy we described in Table 1, for both fixed and reshuffled splits. We always use 80/20 train-validation splits for holdout and 5-fold CVs, so that training set size (and negative estimation bias) are the same. Anytime test performance of an HPO run is assessed by re-training the current incumbent (i.e. the best HPC until the current HPO iteration based on validation performance) on all available train and validation data and evaluating its performance on the outer test set. Note we do this for scientific evaluation in this experiment; obviously, this is not possible in practice. Using random search allows us to record various metrics and afterwards simulate optimizing for different ones, specifically, we recorded accuracy, area under the ROC curve (ROC AUC) and logloss.
@@
We also investigated the effect of reshuffling on two state-of-the-art BO variants (Eggensperger et al., 2021; Turner et al., 2021), namely HEBO (Cowen-Rivers et al., 2022) and SMAC3 (Lindauer et al., 2022). The experimental design was the same as for random search, except for the budget, which we reduced from 500 HPCs to 250 HPCs, and only optimized ROC AUC.